Record the following sequential entries for a company which begins the year with 50,000 shares of $1 par common stock outstanding. For items that say "declare and issue", record both transactions (declaration and issuance) separately. If no entry is required, write NA.

a.) Declare and issue a 30% stock dividend. The market price is $32

b.) Declare and issue a 10% stock dividend. The market price is $22

c.) Declare and issue a 2 for 1 stock split

d.) Declare and issue a property divididend (cost = $50,000; FMV = 70,000)

e.) Declare and issue a liquidating dividend of $20,000

f.) After all of the above how many shares are outstandng?
Answer #1

Part A

declaration of stock dividend

Account titles and explanation Debit Credit
Retained earnings (50000*30%*32) 480000
Common stock dividend distributable (50000*30%*1) 15000
paid in capital in excess of par (50000*30%*31) 465000

Issuance of stock dividend

Account titles and explanation Debit Credit
Common stock dividend distributable (50000*30%*1) 15000
common stock 15000

Part B

declaration of stock dividend

Account titles and explanation Debit Credit
Retained earnings ((50000+(50000*30%))*10%*22) 143000
Common stock dividend distributable ((50000+(50000*30%))*10%*1) 6500
paid in capital in excess of par (50000*30%*31) 136500

Issuance of stock dividend

Account titles and explanation Debit Credit
Common stock dividend distributable ((50000+(50000*30%))*10%*1) 6500
common stock 6500

Part C

NA

There is no journal entry in case of stock split however the number of outstanding shares changes. total number of outstanding shares before stock split = 50000+15000+6500 = 71500

after stock split total number of outstanding share will be = 71500*2 = 143000

Part D

Declaration of property dividend

Account titles and explanation Debit Credit
Retained earnings 70000
Property Dividends Payable 70000

Issuance of property dividend

Account titles and explanation Debit Credit
Property Dividends Payable 70000
Property (asset distributed) 70000

Part E

there will no enty during the declaration of liquidating divided

issuance of liquidating dividend

Account titles and explanation Debit Credit
Retained earnings 20000
Property Dividends Payable 20000

Part F

After all of the above 143000 shares are outstanding because property dividend and liquidating dividend have not effect on outstanding shares.